OpenStudy (brucebaner):

What is the value of the expression? A. 5−20 B. 5−12 C. 54 D. 5-4

OpenStudy (brucebaner):

@AriPotta

OpenStudy (brucebaner):

plz help me @AriPotta

OpenStudy (brucebaner):

@Jhannybean

OpenStudy (brucebaner):

can you help me @calculusxy

OpenStudy (calculusxy):

625

OpenStudy (brucebaner):

what does 625 mean @calculusxy

OpenStudy (brucebaner):

can you help me @perl

OpenStudy (perl):

my draw button does not work

OpenStudy (perl):

5^(-6) = 1/5^6

OpenStudy (brucebaner):

5^(-6) = 1/5^6 @perl mean

OpenStudy (perl):

thats right

OpenStudy (perl):

1 / ( 5*5*5*5*5*5)

OpenStudy (perl):

brb, let me reboot my openstudy

OpenStudy (brucebaner):

so what letter is it

OpenStudy (perl):

sorry i can't give answer , please see code of conduct

OpenStudy (brucebaner):

you @Michele_Laino

OpenStudy (brucebaner):

can you help me @Michele_Laino

OpenStudy (brucebaner):

help me plz @Michele_Laino and @Jhannybean

OpenStudy (michele_laino):

I can write this: \[\frac{ 5^{-6} *5^{2}}{ 5^{-8} }=\frac{ 5^{-4} }{ 5^{-8} }=5^{-4}*5^{8}=5^{4}=\] =625

OpenStudy (jhannybean):

Your answer choices do not make any sense to me.

OpenStudy (brucebaner):

so what letter is 625

OpenStudy (michele_laino):

maybe there is a printing error

OpenStudy (michele_laino):

I think that 54 is a Printing error, I think that author wanted to mean 5^4

OpenStudy (michele_laino):

because 54 is not divisible by 5 or powers of 5

OpenStudy (brucebaner):

its C

OpenStudy (michele_laino):

yes I think!

OpenStudy (brucebaner):

can you help me on in a another question

OpenStudy (michele_laino):

ok!